# SynBullying: A Multi-LLM Synthetic Conversational Dataset for Cyberbullying Detection

### v2

@@ 1. Introduction
Cyberbullying (CB) is a pervasive form of online aggression that disproportionately affects children and adolescents Hinduja and Patchin (2014). It is inherently interactional and context-dependent, emerging through multi-turn exchanges among bullies, victims, and bystanders Sheth et al. (2022). While context is central to most abusive language detection tasks, CB presents unique challenges because the harmfulness of a message often depends on conversational history, relational cues, and patterns of escalation, rather than being apparent from a single message.Ziems et al. (2020). Numerous datasets have been developed for the detection of CB and toxicity Van Hee et al. (2018); Chatzakou et al. (2017); Wulczyn et al. (2017). Although these corpora have advanced abuse and harassment classification, most are post-level or message-level datasets, where annotations are independent of conversational context Ziems et al. (2020). Consequently, models trained on them often fail to capture the interactional intent, indirect aggression, and the discourse framing that define bullying discourse. Moreover, adolescent online communication is linguistically dynamic, shaped by evolving slang, emojis, multimodal references, and coded expressions Dembe (2024); McGillivray et al. (2022). Models trained on static data struggle to generalize to emerging or obfuscated bullying patterns. Collecting authentic conversational data from minors is constrained by ethical, legal, and psychological considerations Facca et al. (2020). Annotation of harmful content is resource-intensive and ethically sensitive, potentially exposing annotators to distress and re-traumatization AlEmadi and Zaghouani (2024). These barriers have created a data scarcity problem, limiting progress in context-aware CB detection and computational modeling of peer aggression. To address these limitations, we introduce SynBullying, a synthetic multi-LLM conversational dataset for studying and detecting CB. SynBullying provides a viable, scalable and ethically safe alternative to human data collection by leveraging large language models (LLMs) to simulate realistic bullying interactions. The dataset offers (i) conversational structure, capturing multi-turn exchanges rather than isolated posts; (ii) context-aware annotations, where harmfulness is assessed within the conversational flow considering context, intent, and discourse dynamics; and (iii) fine-grained labeling, covering various CB categories for detailed linguistic and behavioral analysis. We evaluate SynBullying across six dimensions and examine its utility by testing its performance as standalone training data and as an augmentation source for CB classification.
@@ 3.1. Authentic Dataset
To evaluate the realism of our synthetic CB data, we benchmark it against an authentic CB dataset created through teen role-play sessions Sprugnoli et al. (2018). We use the English version of this dataset Verma et al. (2023). Each simulated conversation assigns participants to predefined roles that mirror real-world CB dynamics: Victim, Bully, Bully Supporter, and Victim Supporter. This structured setup enables the emergence of socially meaningful power hierarchies commonly observed in CB interactions. All conversations are initiated using one of four pre-designed CB-triggering scenarios (labeled A–D), each describing a realistic situation that could lead to peer conflict and harassment. All messages in the dataset were manually annotated by expert annotators using the fine-grained CB types introduced by Van Hee et al. (2015)111Since we only annotate harmful messages with a CB types in this paper, the “Defense” category is excluded.: Threat or Blackmail, Insult General, Insult Body Shame, Insult Discrimination Sexism, Insult Discrimination Racism, Insult Attacking Relatives, Curse or Exclusion, Defamation, Sexual Talk Harmless, Sexual Talk Harassment, Encouragement to Harassment, Other.
@@ 3.2.1. Synthetic Data Generation
We target the generation of multi-turn conversations that explicitly contain CB. To this end, we define a role-based prompt in which eleven fictional teenage participants are assigned specific roles: one victim, two bullies, four victim supporters, and four bully supporters. To encourage the model to produce harmful content within a safe and research-driven setting, we frame the task as academic work conducted for CB detection. The model is then provided with a predefined CB scenario and is instructed to generate realistic, profanity-rich CB conversations. Models occasionally refuse to produce harmful content; in such cases, we reissue the prompt until the desired number of conversations is obtained. To ensure consistency with the authentic dataset, we embed the original role-play scenarios (A–D) from Sprugnoli et al. (2018), which guide the narrative progression and participant interactions.222The final version of this paper will show the prompts used for synthetic data generation in an appendix. Finally, the LLMs generate cyberbullying (CB) conversations as sequences of ordered messages exchanged among participants. Each message is assigned a specific role (Victim, Bully 1, Bully 2, Victim Supporter 1–4, or Bully Supporter 1–4) reflecting the social dynamics of the interaction. Collectively, the messages in each conversation depict a coherent CB incident occurring between these roles.
@@ 5.5. CB Type Distributions
Figure 4 presents the distribution of CB types across datasets333Since a single message may correspond to multiple CB types, the sum of percentages can exceed 100%.. Table 8 reports JSD and p-values for CB-type distributions. Insult_General is the most frequent type in all datasets. Secondary CB types vary across LLMs. GPT-4o best preserves secondary types such as Curse_or_Exclusion and Defamation, whereas Grok and LLaMA overrepresent discrimination-related insults. GPT-4o exhibits the closest match to the authentic CB-type distribution (JSD = 0.0317), while Grok and LLaMA deviate more. Although GPT-4o is closest, all synthetic CB-type distributions are significantly different from authentic data (p 1e-28), reflecting model-specific biases in generating harmful content.
@@ 9. Bibliographical References
- AlEmadi and Zaghouani (2024) Maryam M AlEmadi and Wajdi Zaghouani. 2024. Emotional toll and coping strategies: Navigating the effects of annotating hate speech data. In Proceedings of the Workshop on Legal and Ethical Issues in Human Language Technologies@ LREC-COLING 2024, pages 66–72.
- Anaby-Tavor et al. (2020) Ateret Anaby-Tavor, Boaz Carmeli, Esther Goldbraich, Amir Kantor, George Kour, Segev Shlomov, Naama Tepper, and Naama Zwerdling. 2020. Do not have enough data? deep learning to the rescue! Proceedings of the AAAI Conference on Artificial Intelligence, 34(05):7383–7390.
- Bauman (2015) Sheri Bauman. 2015. Types of cyberbullying. Cyberbullying: What counselors need to know, pages 53–58.
@@
- Bojić et al. (2025) Ljubiša Bojić, Olga Zagovora, Asta Zelenkauskaite, Vuk Vuković, Milan Čabarkapa, Selma Veseljević Jerković, and Ana Jovančević. 2025. Comparing large language models and human annotators in latent content analysis of sentiment, political leaning, emotional intensity and sarcasm. Scientific reports, 15(1):11477.
@@
- Bonifacio et al. (2022) Luiz Bonifacio, Hugo Abonizio, Marzieh Fadaee, and Rodrigo Nogueira. 2022. InPars: Unsupervised dataset generation for information retrieval. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR ’22, pages 2387–2392, New York, NY, USA. Association for Computing Machinery.
- Chatzakou et al. (2017) Despoina Chatzakou, Nicolas Kourtellis, Jeremy Blackburn, Emiliano De Cristofaro, Gianluca Stringhini, and Athena Vakali. 2017. Mean birds: Detecting aggression and bullying on twitter. In Proceedings of the 2017 ACM on web science conference, pages 13–22.
@@
- Dembe (2024) Tony Dembe. 2024. The impact of social media on language evolution. European Journal of Linguistics, 3(3):1–14.
@@
- Devlin et al. (2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019. BERT: Pre-training of deep bidirectional transformers for language understanding. In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers), pages 4171–4186, Minneapolis, Minnesota. Association for Computational Linguistics.
@@
- Ejaz et al. (2024) Naveed Ejaz, Fakhra Kashif, and Salimur Choudhury. 2024. A multi-faceted semi-synthetic dataset for automated cyberbullying detection.
@@
- Emmery et al. (2021) Chris Emmery, Ben Verhoeven, Guy De Pauw, Gilles Jacobs, Cynthia Van Hee, Els Lefever, Bart Desmet, Véronique Hoste, and Walter Daelemans. 2021. Current limitations in cyberbullying detection: On evaluation criteria, reproducibility, and data scarcity. Language Resources and Evaluation, 55:597–633.
@@
- Facca et al. (2020) Danica Facca, Maxwell J Smith, Jacob Shelley, Daniel Lizotte, and Lorie Donelle. 2020. Exploring the ethical issues in research using digital data collection strategies with minors: A scoping review. Plos one, 15(8):e0237875.
@@
- Ghanadian et al. (2024) Hamideh Ghanadian, Isar Nejadgholi, and Hussein Al Osman. 2024. Socially aware synthetic data generation for suicidal ideation detection using large language models. IEEE Access, 12:14350–14363.
@@
- He et al. (2022) Xuanli He, Islam Nassar, Jamie Kiros, Gholamreza Haffari, and Mohammad Norouzi. 2022. Generate, annotate, and learn: NLP with synthetic text. Transactions of the Association for Computational Linguistics, 10:826–842.
@@
- He et al. (2021) Xuanli He, Islam Nassar, Jamie Ryan Kiros, Gholamreza Haffari, and Mohammad Norouzi. 2021. Generate, annotate, and learn: Generative models advance self-training and knowledge distillation. Manuscript submitted to ICLR 2022.
@@
- Hinduja and Patchin (2014) Sameer Hinduja and Justin W Patchin. 2014. Bullying beyond the schoolyard: Preventing and responding to cyberbullying. Corwin press.
@@
- Hui et al. (2024) Zheng Hui, Zhaoxiao Guo, Hang Zhao, Juanyong Duan, and Congrui Huang. 2024. ToxiCraft: A novel framework for synthetic generation of harmful information. In Findings of the Association for Computational Linguistics: EMNLP 2024, pages 16632–16647, Miami, Florida, USA. Association for Computational Linguistics.
@@
- Hutto and Gilbert (2014) C.J. Hutto and Eric Gilbert. 2014. VADER: A parsimonious rule-based model for sentiment analysis of social media text. Proceedings of the Eighth International Conference on Weblogs and Social Media (ICWSM-14), 8(1):216–225.
@@
- Kazemi et al. (2025) Arefeh Kazemi, Sri Balaaji Natarajan Kalaivendan, Joachim Wagner, Hamza Qadeer, Kanishk Verma, and Brian Davis. 2025. Synthetic vs. gold: The role of llm generated labels and data in cyberbullying detection. In Proceedings of the 15th International Conference on Recent Advances in Natural Language Processing - Natural Language Processing in the Generative AI era, pages 531–540, Varna, Bulgaria. INCOMA Ltd., Shoumen, Bulgaria.
- Kumar et al. (2024) Yulia Kumar, Kuan Huang, Angelo Perez, Guohao Yang, J. Jenny Li, Patricia Morreale, Dov Kruger, and Raymond Jiang. 2024. Bias and cyberbullying detection and data generation using transformer artificial intelligence models and top large language models. Electronics, 13(17).
- Landis and Koch (1977) J Richard Landis and Gary G Koch. 1977. The measurement of observer agreement for categorical data. biometrics, pages 159–174.
@@
- Leung et al. (2018) Angel NM Leung, Natalie Wong, and JoAnn M Farver. 2018. You are what you read: The belief systems of cyber-bystanders on social networking sites. Frontiers in psychology, 9:365.
@@
- Loper and Bird (2002) Edward Loper and Steven Bird. 2002. NLTK: The natural language toolkit. In Proceedings of the ACL-02 Workshop on Effective Tools and Methodologies for Teaching Natural Language Processing and Computational Linguistics, pages 63–70, Philadelphia, Pennsylvania, USA. Association for Computational Linguistics.
@@
- McCarthy and Jarvis (2010) Philip M McCarthy and Scott Jarvis. 2010. Mtld, vocd-d, and hd-d: A validation study of sophisticated approaches to lexical diversity assessment. Behavior research methods, 42(2):381–392.
@@
- McGillivray et al. (2022) Barbara McGillivray, Malithi Alahapperuma, Jonathan Cook, Chiara Di Bonaventura, Albert Merono Penuela, Gareth Tyson, and Steven Wilson. 2022. Leveraging time-dependent lexical features for offensive language detection. In Proceedings of the 1st Workshop of Ever Evolving NLP, EMNLP 2022.
@@
- Meng et al. (2022) Yu Meng, Jiaxin Huang, Yu Zhang, and Jiawei Han. 2022. Generating training data with language models: Towards zero-shot language understanding. In Advances in Neural Information Processing Systems, volume 35, pages 462–477. Curran Associates, Inc.
@@
- Meta (2024) Meta. 2024. llama-models/models/llama3_3. Git repository of GitHub user “meta-llama”.
- Nadali et al. (2013) Samaneh Nadali, Masrah Azrifah Azmi Murad, Nurfadhlina Mohamad Sharef, Aida Mustapha, and Somayeh Shojaee. 2013. A review of cyberbullying detection: An overview. In 2013 13th International Conference on Intellient Systems Design and Applications, pages 325–330. IEEE.
@@
- Ollagnier et al. (2022) Anaïs Ollagnier, Elena Cabrio, Serena Villata, and Catherine Blaya. 2022. CyberAgressionAdo-v1: a dataset of annotated online aggressions in French collected through a role-playing game. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 867–875, Marseille, France. European Language Resources Association.
- OpenAI (2024) OpenAI. 2024. GPT-4o system card.
@@
- Patchin and Hinduja (2006) Justin W Patchin and Sameer Hinduja. 2006. Bullies move beyond the schoolyard: A preliminary look at cyberbullying. Youth Violence and Juvenile Justice, 4(2):148–169.
@@
- Pérez et al. (2024) Jaime Pérez, Mario Castro, Edmond Awad, and Gregorio López. 2024. Generation of probabilistic synthetic data for serious games: A case study on cyberbullying. Knowledge-Based Systems, 286:111440.
@@
- Sason (2022) Igal Sason. 2022. Divergence measures: mathematical foundations and applications in information-theoretic and statistical problems.
@@
- Schmidhuber and Kruschwitz (2024) Maximilian Schmidhuber and Udo Kruschwitz. 2024. LLM-based synthetic datasets: Applications and limitations in toxicity detection. In Proceedings of the Fourth Workshop on Threat, Aggression & Cyberbullying @ LREC-COLING-2024, pages 37–51, Torino, Italia. ELRA and ICCL.
@@
- Sheth et al. (2022) Amit Sheth, Valerie L Shalin, and Ugur Kursuncu. 2022. Defining and detecting toxicity on social media: context and knowledge are key. Neurocomputing, 490:312–318.
@@
- Slonje et al. (2013) Robert Slonje, Peter K Smith, and Ann Frisén. 2013. The nature of cyberbullying, and strategies for prevention. Computers in human behavior, 29(1):26–32.
@@
- Song and Oh (2018) Jiyeon Song and Insoo Oh. 2018. Factors influencing bystanders’ behavioral reactions in cyberbullying situations. Computers in Human Behavior, 78:273–282.
- Sprugnoli et al. (2018) Rachele Sprugnoli, Stefano Menini, Sara Tonelli, Filippo Oncini, and Enrico Piras. 2018. Creating a WhatsApp dataset to study pre-teen cyberbullying. In Proceedings of the 2nd Workshop on Abusive Language Online (ALW2), pages 51–59, Brussels, Belgium. Association for Computational Linguistics.
@@
- Tari et al. (2025) Henry Tari, Nojus Sereiva, Rishabh Kaushal, Thales Bertaglia, and Adriana Iamnitchi. 2025. Towards high-fidelity synthetic multi-platform social media datasets via large language models.
@@
- Team (2020) Unitary Team. 2020. Toxicbert. https://huggingface.co/unitary/toxic-bert.
- Van Hee et al. (2018) Cynthia Van Hee, Gilles Jacobs, Chris Emmery, Bart Desmet, Els Lefever, Ben Verhoeven, Guy De Pauw, Walter Daelemans, and Véronique Hoste. 2018. Automatic detection of cyberbullying in social media text. PloS one, 13(10):e0203794.
- Van Hee et al. (2015) Cynthia Van Hee, Ben Verhoeven, Els Lefever, Guy De Pauw, Véronique Hoste, and Walter Daelemans. 2015. Guidelines for the fine-grained analysis of cyberbullying, version 1.0. T3 Technical Report – LT3 15-01.
@@
- Verma et al. (2023) Kanishk Verma, Maja Popović, Alexandros Poulis, Yelena Cherkasova, Cathal Ó hÓbáin, Angela Mazzone, Tijana Milosevic, and Brian Davis. 2023. Leveraging machine translation for cross-lingual fine-grained cyberbullying classification amongst pre-adolescents. Natural Language Engineering, 29(6):1458–1480.
@@
- Wang et al. (2024) Junda Wang, Zonghai Yao, Zhichao Yang, Huixue Zhou, Rumeng Li, Xun Wang, Yucheng Xu, and Hong Yu. 2024. NoteChat: A dataset of synthetic patient-physician conversations conditioned on clinical notes. In Findings of the Association for Computational Linguistics: ACL 2024, pages 15183–15201, Bangkok, Thailand. Association for Computational Linguistics.
@@
- Wulczyn et al. (2017) Ellery Wulczyn, Nithum Thain, and Lucas Dixon. 2017. Ex machina: Personal attacks seen at scale. In Proceedings of the 26th international conference on world wide web, pages 1391–1399.
@@
- xAI (2024) xAI. 2024. Bringing Grok to everyone. Product news.
@@
- Xu et al. (2012) Jun-Ming Xu, Kwang-Sung Jun, Xiaojin Zhu, and Amy Bellmore. 2012. Learning from bullying traces in social media. In Proceedings of the 2012 conference of the North American chapter of the association for computational linguistics: Human language technologies, pages 656–666.
@@
- Yoo et al. (2021) Kang Min Yoo, Dongju Park, Jaewook Kang, Sang-Woo Lee, and Woomyoung Park. 2021. GPT3Mix: Leveraging large-scale language models for text augmentation. In Findings of the Association for Computational Linguistics: EMNLP 2021, pages 2225–2239, Punta Cana, Dominican Republic. Association for Computational Linguistics.

### v3

@@ 1. Introduction
Cyberbullying (CB) is a pervasive form of online aggression that disproportionately affects children and adolescents Hinduja and Patchin (2014). It is inherently interactional and context-dependent, emerging through multi-turn exchanges among bullies, victims, and bystanders Sheth et al. (2022). While context is central to most abusive language detection tasks, CB presents unique challenges because the harmfulness of a message often depends on conversational history, relational cues, and patterns of escalation, rather than being apparent from a single message.Ziems et al. (2020). Numerous datasets have been developed for the detection of CB and toxicity Van Hee et al. (2018); Chatzakou et al. (2017); Wulczyn et al. (2017). Although these corpora have advanced abuse and harassment classification, most are post-level or message-level datasets, where annotations are independent of conversational context Ziems et al. (2020). Consequently, models trained on them often fail to capture the interactional intent, indirect aggression, and the discourse framing that define bullying discourse. Moreover, adolescent online communication is linguistically dynamic, shaped by evolving slang, emojis, multimodal references, and coded expressions Dembe (2024); McGillivray et al. (2022). Models trained on static data struggle to generalize to emerging or obfuscated bullying patterns. Collecting authentic conversational data from minors is constrained by ethical, legal, and psychological considerations Facca et al. (2020). Annotation of harmful content is resource-intensive and ethically sensitive, potentially exposing annotators to distress and re-traumatization AlEmadi and Zaghouani (2024). These barriers have created a data scarcity problem, limiting progress in context-aware CB detection and computational modeling of peer aggression. To address these limitations, we introduce SynBullying, a synthetic multi-LLM conversational dataset for studying and detecting CB.111Publicly available at https://huggingface.co/datasets/arrkaa-NLP/SynBullying. SynBullying provides a viable, scalable and ethically safe alternative to human data collection by leveraging large language models (LLMs) to simulate realistic bullying interactions. The dataset offers (i) conversational structure, capturing multi-turn exchanges rather than isolated posts; (ii) context-aware annotations, where harmfulness is assessed within the conversational flow considering context, intent, and discourse dynamics; and (iii) fine-grained labeling, covering various CB categories for detailed linguistic and behavioral analysis. We evaluate SynBullying across six dimensions and examine its utility by testing its performance as standalone training data and as an augmentation source for CB classification.
@@ 3.1. Authentic Dataset
To evaluate the realism of our synthetic CB data, we benchmark it against an authentic CB dataset created through teen role-play sessions Sprugnoli et al. (2018). We use the English version of this dataset Verma et al. (2023). Each simulated conversation assigns participants to predefined roles that mirror real-world CB dynamics: Victim, Bully, Bully Supporter, and Victim Supporter. This structured setup enables the emergence of socially meaningful power hierarchies commonly observed in CB interactions. All conversations are initiated using one of four pre-designed CB-triggering scenarios (labeled A–D), each describing a realistic situation that could lead to peer conflict and harassment. All messages in the dataset were manually annotated by expert annotators using the fine-grained CB types introduced by Van Hee et al. (2015)222Since we only annotate harmful messages with a CB types in this paper, the “Defense” category is excluded.: Threat or Blackmail, Insult General, Insult Body Shame, Insult Discrimination Sexism, Insult Discrimination Racism, Insult Attacking Relatives, Curse or Exclusion, Defamation, Sexual Talk Harmless, Sexual Talk Harassment, Encouragement to Harassment, Other.
@@ 3.2.1. Synthetic Data Generation
We target the generation of multi-turn conversations that explicitly contain CB. To this end, we define a role-based prompt in which eleven fictional teenage participants are assigned specific roles: one victim, two bullies, four victim supporters, and four bully supporters. To encourage the model to produce harmful content within a safe and research-driven setting, we frame the task as academic work conducted for CB detection. The model is then provided with a predefined CB scenario and is instructed to generate realistic, profanity-rich CB conversations. Models occasionally refuse to produce harmful content; in such cases, we reissue the prompt until the desired number of conversations is obtained. To ensure consistency with the authentic dataset, we embed the original role-play scenarios (A–D) from Sprugnoli et al. (2018), which guide the narrative progression and participant interactions.333The final version of this paper will show the prompts used for synthetic data generation in an appendix. Finally, the LLMs generate cyberbullying (CB) conversations as sequences of ordered messages exchanged among participants. Each message is assigned a specific role (Victim, Bully 1, Bully 2, Victim Supporter 1–4, or Bully Supporter 1–4) reflecting the social dynamics of the interaction. Collectively, the messages in each conversation depict a coherent CB incident occurring between these roles.
@@ 5.5. CB Type Distributions
Figure 4 presents the distribution of CB types across datasets444Since a single message may correspond to multiple CB types, the sum of percentages can exceed 100%.. Table 8 reports JSD and p-values for CB-type distributions. Insult_General is the most frequent type in all datasets. Secondary CB types vary across LLMs. GPT-4o best preserves secondary types such as Curse_or_Exclusion and Defamation, whereas Grok and LLaMA overrepresent discrimination-related insults. GPT-4o exhibits the closest match to the authentic CB-type distribution (JSD = 0.0317), while Grok and LLaMA deviate more. Although GPT-4o is closest, all synthetic CB-type distributions are significantly different from authentic data (p 1e-28), reflecting model-specific biases in generating harmful content.
@@ 8. Ethical Considerations
9. Acknowledgements
This research is supported by the Disruptive Technologies Innovation Fund (DTIF) under the project “Cilter: Protecting Children Online” Grant No. DT 2021 0362 from the Department of Enterprise, Trade and Employment in Ireland and administered by Enterprise Ireland (EI). This research was conducted with the financial support of Science Foundation Ireland under Grant Agreement No. 13/RC/2106_P2 at the ADAPT SFI Research Centre at Dublin City University. ADAPT, the SFI Research Centre for AI-Driven Digital Content Technology, is funded by Science Foundation Ireland through the SFI Research Centres Programme. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
@@ 10. Bibliographical References
- Emotional toll and coping strategies: navigating the effects of annotating hate speech data. In Proceedings of the Workshop on Legal and Ethical Issues in Human Language Technologies@ LREC-COLING 2024, pp. 66–72. Cited by: §1.
- Do not have enough data? deep learning to the rescue!. Proceedings of the AAAI Conference on Artificial Intelligence 34 (05), pp. 7383–7390. External Links: Link, Document Cited by: §2.
- Types of cyberbullying. Cyberbullying: What counselors need to know, pp. 53–58. Cited by: §2.
@@
- Comparing large language models and human annotators in latent content analysis of sentiment, political leaning, emotional intensity and sarcasm. Scientific reports 15 (1), pp. 11477. Cited by: §4.3.
@@
- InPars: unsupervised dataset generation for information retrieval. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR ’22, New York, NY, USA, pp. 2387–2392. External Links: ISBN 9781450387323, Link, Document Cited by: §2.
- Mean birds: detecting aggression and bullying on twitter. In Proceedings of the 2017 ACM on web science conference, pp. 13–22. Cited by: §1.
@@
- The impact of social media on language evolution. European Journal of Linguistics 3 (3), pp. 1–14. Cited by: §1.
@@
- BERT: pre-training of deep bidirectional transformers for language understanding. In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers), J. Burstein, C. Doran, and T. Solorio (Eds.), Minneapolis, Minnesota, pp. 4171–4186. External Links: Link, Document Cited by: §5.6.1.
@@
- A multi-faceted semi-synthetic dataset for automated cyberbullying detection. External Links: 2402.10231, Link Cited by: §2.
@@
- Current limitations in cyberbullying detection: on evaluation criteria, reproducibility, and data scarcity. Language Resources and Evaluation 55, pp. 597–633. Cited by: §2.
@@
- Exploring the ethical issues in research using digital data collection strategies with minors: a scoping review. Plos one 15 (8), pp. e0237875. Cited by: §1.
@@
- Socially aware synthetic data generation for suicidal ideation detection using large language models. IEEE Access 12 (), pp. 14350–14363. External Links: Document Cited by: §2.
@@
- Generate, annotate, and learn: NLP with synthetic text. Transactions of the Association for Computational Linguistics 10, pp. 826–842. External Links: Link, Document Cited by: §2.
@@
- Generate, annotate, and learn: generative models advance self-training and knowledge distillation. Note: manuscript submitted to ICLR 2022 External Links: Link Cited by: §2.
@@
- Bullying beyond the schoolyard: preventing and responding to cyberbullying. Corwin press. Cited by: §1.
@@
- ToxiCraft: a novel framework for synthetic generation of harmful information. In Findings of the Association for Computational Linguistics: EMNLP 2024, Y. Al-Onaizan, M. Bansal, and Y. Chen (Eds.), Miami, Florida, USA, pp. 16632–16647. External Links: Link, Document Cited by: §2.
@@
- VADER: a parsimonious rule-based model for sentiment analysis of social media text. Proceedings of the Eighth International Conference on Weblogs and Social Media (ICWSM-14) 8 (1), pp. 216–225. External Links: Link, Document Cited by: §5.2.
@@
- Synthetic vs. gold: the role of llm generated labels and data in cyberbullying detection. In Proceedings of the 15th International Conference on Recent Advances in Natural Language Processing - Natural Language Processing in the Generative AI era, Varna, Bulgaria, pp. 531–540. External Links: Link Cited by: §A.1, §A.1, §2.
- Bias and cyberbullying detection and data generation using transformer artificial intelligence models and top large language models. Electronics 13 (17). External Links: Link, ISSN 2079-9292, Document Cited by: §2.
- The measurement of observer agreement for categorical data. biometrics, pp. 159–174. Cited by: §4.1.
@@
- You are what you read: the belief systems of cyber-bystanders on social networking sites. Frontiers in psychology 9, pp. 365. Cited by: §2.
@@
- NLTK: the natural language toolkit. In Proceedings of the ACL-02 Workshop on Effective Tools and Methodologies for Teaching Natural Language Processing and Computational Linguistics, Philadelphia, Pennsylvania, USA, pp. 63–70. External Links: Link, Document Cited by: §5.2.
@@
- MTLD, vocd-d, and hd-d: a validation study of sophisticated approaches to lexical diversity assessment. Behavior research methods 42 (2), pp. 381–392. Cited by: §5.1.
@@
- Leveraging time-dependent lexical features for offensive language detection. In Proceedings of the 1st Workshop of Ever Evolving NLP, EMNLP 2022, Cited by: §1.
@@
- Generating training data with language models: towards zero-shot language understanding. In Advances in Neural Information Processing Systems, S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Eds.), Vol. 35, pp. 462–477. External Links: Link Cited by: §2.
@@
- Llama-models/models/llama3_3. Note: Git repository of GitHub user “meta-llama” External Links: Link Cited by: §3.2.
- A review of cyberbullying detection: an overview. In 2013 13th International Conference on Intellient Systems Design and Applications, pp. 325–330. Cited by: §2.
@@
- CyberAgressionAdo-v1: a dataset of annotated online aggressions in French collected through a role-playing game. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, Marseille, France, pp. 867–875. External Links: Link Cited by: §2.
- GPT-4o system card. External Links: Document, 2410.21276v1, Link Cited by: §3.2.
@@
- Bullies move beyond the schoolyard: a preliminary look at cyberbullying. Youth Violence and Juvenile Justice 4 (2), pp. 148–169. External Links: Document Cited by: §2.
@@
- Generation of probabilistic synthetic data for serious games: a case study on cyberbullying. Knowledge-Based Systems 286, pp. 111440. External Links: Document Cited by: §2.
@@
- Divergence measures: mathematical foundations and applications in information-theoretic and statistical problems. Vol. 24, MDPI. Cited by: §5.4.
@@
- LLM-based synthetic datasets: applications and limitations in toxicity detection. In Proceedings of the Fourth Workshop on Threat, Aggression & Cyberbullying @ LREC-COLING-2024, R. Kumar, A. Kr. Ojha, S. Malmasi, B. R. Chakravarthi, B. Lahiri, S. Singh, and S. Ratan (Eds.), Torino, Italia, pp. 37–51. External Links: Link Cited by: §2.
@@
- Defining and detecting toxicity on social media: context and knowledge are key. Neurocomputing 490, pp. 312–318. Cited by: §1.
@@
- The nature of cyberbullying, and strategies for prevention. Computers in human behavior 29 (1), pp. 26–32. Cited by: §2.
@@
- Factors influencing bystanders’ behavioral reactions in cyberbullying situations. Computers in Human Behavior 78, pp. 273–282. Cited by: §2.
- Creating a WhatsApp dataset to study pre-teen cyberbullying. In Proceedings of the 2nd Workshop on Abusive Language Online (ALW2), D. Fišer, R. Huang, V. Prabhakaran, R. Voigt, Z. Waseem, and J. Wernimont (Eds.), Brussels, Belgium, pp. 51–59. External Links: Link, Document Cited by: §A.1, §2, §3.1, §3.2.1.
@@
- Towards high-fidelity synthetic multi-platform social media datasets via large language models. External Links: 2505.02858, Link Cited by: §2.
@@
- ToxicBERT. Note: https://huggingface.co/unitary/toxic-bert Cited by: §5.2.
- Automatic detection of cyberbullying in social media text. PloS one 13 (10), pp. e0203794. Cited by: §1, §2.
- Guidelines for the fine-grained analysis of cyberbullying, version 1.0. Note: T3 Technical Report – LT3 15-01 External Links: Link Cited by: §3.1, §4.1, §4.2.
@@
- Leveraging machine translation for cross-lingual fine-grained cyberbullying classification amongst pre-adolescents. Natural Language Engineering 29 (6), pp. 1458–1480. External Links: Document Cited by: §3.1.
@@
- NoteChat: a dataset of synthetic patient-physician conversations conditioned on clinical notes. In Findings of the Association for Computational Linguistics: ACL 2024, L. Ku, A. Martins, and V. Srikumar (Eds.), Bangkok, Thailand, pp. 15183–15201. External Links: Link, Document Cited by: §2.
@@
- Ex machina: personal attacks seen at scale. In Proceedings of the 26th international conference on world wide web, pp. 1391–1399. Cited by: §1.
@@
- Bringing Grok to everyone. Note: Product news External Links: Link Cited by: §3.2.
@@
- Learning from bullying traces in social media. In Proceedings of the 2012 conference of the North American chapter of the association for computational linguistics: Human language technologies, pp. 656–666. Cited by: §4.2.
@@
- GPT3Mix: leveraging large-scale language models for text augmentation. In Findings of the Association for Computational Linguistics: EMNLP 2021, M. Moens, X. Huang, L. Specia, and S. W. Yih (Eds.), Punta Cana, Dominican Republic, pp. 2225–2239. External Links: Link, Document Cited by: §2.